OpenStudy (anonymous):

find the domain and range of f(x) = 1/(x^2-1) please show me the steps to find the range.

OpenStudy (agent0smith):

One way to find the range is to find the inverse function. The *domain* of the inverse is the *range* of the original. Domain is easier to find than range.

OpenStudy (agent0smith):

\[\Large y = \frac{1} {x^2-1}\] \[\Large x^2-1 = \frac{1} {y}\] \[\Large x =\pm \sqrt { \frac{1} {y}+1}\] It will be where \[\large \frac{ 1 }{ y} +1 \ge 0\]

OpenStudy (anonymous):

i want to understand how you write 1/y+1 >= 0 in the interval notation

OpenStudy (agent0smith):

i can't find an easy way to show it algebraically...\[\large \frac{ 1 }{ y} \ge -1\]but it'll be when y>0 (ie when y is positive 1/y is always >-1) when y=<-1 (remember if y is something like -0.5, 1/-0.5 is -2. But if y is -2, 1/-2 is > -1)